OpenStudy (anonymous):

Compare: five-twelfths to seven-eighths less than greater than equal to None of the above

OpenStudy (anonymous):

less than

OpenStudy (anonymous):

5/12 = 20/48 7/8= 42/48

OpenStudy (anonymous):

20/48 < 42/48

OpenStudy (anonymous):

Explain the steps you would use to compare four-fifteenths and five-sixths. Remember, you can use the forward-slash key (/) to type fractions. For example, type one-fourth as one slash four.